OpenStudy (anonymous):

A package contains 16 candy canes, 6 of which are cracked. If 3 are selected, find the probability of getting exactly two cracked candy canes.

OpenStudy (anonymous):

The probability can be taken here by Combinations. I believe algebra?

OpenStudy (anonymous):

(6 2)*(10 1)/(16 3)

OpenStudy (anonymous):

To randomly select 3 out of 16, of which 2 are crached, you need 6C2*10C1=6!/(4!2!) * 10=Cases in which you've selected 2 cracked candy.

OpenStudy (anonymous):

The next part is the total number of cases. Which is 16C3=(16*15*14)/(3*2)

OpenStudy (anonymous):

You divide the former expression by latter. I hope it helps.